OpenStudy (anonymous):

eliminate the parameter and write the corresponding rectangular equation... PLEASE! : x=4+ 2 cos theta and y=-1+2 sin theta

OpenStudy (anonymous):

is it (x-4)^2/16+(y+1)^2/4=1 ?

OpenStudy (anonymous):

and the graph please

jimthompson5910 (jim_thompson5910):

x = 4 + 2cos(t) x-4 = 2cos(t) (x-4)^2 = (2cos(t))^2 (x-4)^2 = 4cos^2(t) -------------------------- y = -1+2sin(t) y+1 = 2sin(t) (y+1)^2 = (2sin(t))^2 (y+1)^2 = 4sin^2(t) --------------------------------------- So we have (x-4)^2 = 4cos^2(t) and (y+1)^2 = 4sin^2(t) Add the equations to get (x-4)^2 + (y+1)^2 = 4cos^2(t) + 4sin^2(t) (x-4)^2 + (y+1)^2 = 4(cos^2(t) + sin^2(t)) (x-4)^2 + (y+1)^2 = 4(1) (x-4)^2 + (y+1)^2 = 4 So it's a circle center at (4,-1) with a radius of 2

OpenStudy (anonymous):

I got that:) can u show a pic of the graph @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Use a graphing calculator to get

OpenStudy (anonymous):

thats what i got:D

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

I have another one..do mind

jimthompson5910 (jim_thompson5910):

no go ahead

OpenStudy (anonymous):

Ok can you graph these ? the first problem is: evolute of ellipse: x= 2 cos^3 theta and y=4 sin^3 theta.

OpenStudy (anonymous):

And the second problem is : serpentine curve: x=1/2 cot theta and y=4 sin theta cos theta

OpenStudy (anonymous):

are u there @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah just graphing, one sec

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

here's the first graph

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

second is serpentine curve: x=1/2 cot theta and y=4 sin theta cos theta

jimthompson5910 (jim_thompson5910):

and here's the graph for that

OpenStudy (anonymous):

thanks so much!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

hi Jim, how did you plus in cosine to the third into the graphing calculator. Is there some special key to raising cosine theta to the third